OpenStudy (anonymous):

simplify (2/3x^3+1/2x^2-1/4x)-(-1/8x+3/4x^2-1/3x^3)

OpenStudy (espex):

Assuming your problem is this: \[(\frac{2}{3}x^3+\frac{1}{2}x^2 - \frac{1}{4}x)-(-\frac{1}{8}x+\frac{3}{4}x^2-\frac{1}{3}x^3)\] I would suggest you start by distributing the minus sign into the second quantity. Then look at combining like terms, and depending on how far you need to go you could factor them.

OpenStudy (anonymous):

-(x-2) (x+4)/8x^3 -X^2 + 2x – 8/ 8x^3 -x^2 – 2x + 8 / 8x^3

OpenStudy (anonymous):

i still don't get it

OpenStudy (espex):

How far did you get?

OpenStudy (anonymous):

i came up with 1/3^3-1/4x^2-1/8x

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

i got it

OpenStudy (espex):

... Okay, that's great. :)

OpenStudy (anonymous):

\[x^3-1/4^2-1/8x\]

OpenStudy (espex):

Close.. :)

OpenStudy (anonymous):

right....?? whew!

OpenStudy (anonymous):

x^3-1/4x^2-1/8x

OpenStudy (espex):

There ya go.. Although I believe the x is not part of the denominator..

OpenStudy (anonymous):

its not...

OpenStudy (espex):

Now you can look at factoring it if you need to simplify it further.

OpenStudy (anonymous):

thanks so much. now i should do well on my final next week....this was stumping me even in class!!!

OpenStudy (anonymous):

Are you familiar with polynomials?

OpenStudy (espex):

Which order?

OpenStudy (anonymous):

If there is an integer that has a powere of y, that would not be a polynomial correct? such as 3.2x^4+2.4x^y+7.3

OpenStudy (anonymous):

And if on of the number has a power of a fraction that is not a polynomial correct? such as 4x^2+9x^1/3+y^2

OpenStudy (anonymous):

I'm viewing my notes and it seems confusing

OpenStudy (espex):

Correct, as I understand it all exponents must be integers greater than 0.

OpenStudy (espex):

So in your two examples those would NOT be polynomials because 'y' and 1/3 are not integers.

OpenStudy (anonymous):

what if it were ^1/2

OpenStudy (espex):

No, fractions are not integers.

OpenStudy (espex):

exponents can only be 0,1,2,3,... etc.

OpenStudy (anonymous):

can absolute values be polynomials?

OpenStudy (anonymous):

since 1/2 is .50 would that be an exception?

OpenStudy (espex):

No.

OpenStudy (espex):

1/2 is a fraction and thus cannot be a polynomial.

OpenStudy (anonymous):

No? that absoulte values cannot be polynomials?

OpenStudy (espex):

I would argue a conditional on that, since the absolute value is a way of handling a number it would not be a polynomial, but if you look at just the value you might say that x^|5| still has a positive integer as a power so maybe. However by the strict definition of a polynomial, whose exponents can only be 0,1,2,3,4,..etc. then NO, even absolute value would not count.

OpenStudy (anonymous):

thanks